# Argumentatively Coherent Judgmental Forecasting

### v1

@@ 2 Related Work
In Judgmental forecasting, there is some work that tries to incorporate coherence, for example by investigating the best ways of eliciting probabilities from humans Wallsten and Budescu (1995), how incentives and training change forecasters’ abilities Chang et al. (2016), the effect of scoring rules on forecasters Zellner et al. (2021), and how principles help overcome overconfidence Arkes (2001). Some existing solutions Karvetski et al. (2013); Fan et al. (2019); Gibson et al. (2021) show that forecasting accuracy improves by weighting judges’ assessments based on their degree of coherence. Further, Thomson et al. (2019) develop a measure of coherence to identify the degree of independence between sets of forecasts, thus improving the aggregated forecasts. Additionally, Ho et al. (2024) introduce a notion of probabilistic coherence that identifies superior forecasters. Finally, there has been some work on coherence constraints in hierarchical forecasting Rangapuram et al. (2021); Panagiotelis et al. (2023); Rangapuram et al. (2023); Olivares et al. (2023).
@@
Existing research Irwin et al. (2022); Gorur et al. (2023); Toni et al. (2023) explores the integration of computational argumentation for eliciting, representing and evaluating the argumentative reasoning of users. Specifically, Irwin et al. (2022) were the first to apply computational argumentation to support judgmental forecasting. They introduced their notion of coherence based on how forecasters update their predictions and their reasoning behind those updates. However, as mentioned in the introduction, their underlying framework exhibits some limitations in that it can only handle binary questions and does not allow the same arguments to be used for different questions. Further demonstrating the benefits of argumentation in this context is Toni et al. (2023), which incorporates users’ opinions and probabilistic estimates to aid forecasting, but without considering coherence constraints. In addition, the role of argumentation towards forecasting is indirectly advocated by resources such as the FORECAST2023 dataset Gorska et al. (2024), where the rationales from the Hybrid-Forecasting Competition dataset are annotated with support and attack relations.
@@
Recent research demonstrates the potential of LLMs in forecasting, e.g. in time-series forecasting, LLMs have been shown to improve accuracy Jin et al. (2024); Tang et al. (2024); Liu et al. (2024). However, direct use of LLMs in judgmental forecasting has yielded mixed results: while some studies Schoenegger and Park (2023); Abolghasemi et al. (2023) show that using these models directly does not improve forecasting accuracy, others Halawi et al. (2024) have concluded that the current LLMs are nearly as good as human forecasters if they are supported with Retrieval Augmented Generation (RAG). They further point that the aggregation of forecasts from humans and LLMs with RAG achieve the best forecasting accuracy. Combining human insights with LLM-supported advice has also proven effective in accuracy Schoenegger et al. (2024). Our approach in Section 6 incorporates LLMs on forecaster data to improve forecasting accuracy. It does so by adapting, towards forecasting, argLLMs Freedman et al. (2025), an existing system for mining argumentation frameworks from LLMs and use these frameworks to verify claims.
@@ 3 Background
QBAFs are equipped with so-called gradual semantics, assessing the (dialectical) strength of arguments. We use the Discontinuity-Free Quantitative Argumentation Debate (DF-QuAD) gradual semantics Rago et al. (2016) to obtain the strength of the forecasters’ opinions. For a given QBAF $\langle\mathcal{X},\mathcal{A},\mathcal{S},\tau\rangle$, for any $x\in\mathcal{X}$ with $n\geq 0$ attackers with strengths $v_{1},\ldots,v_{n}$, $m\geq 0$ supporters with strengths $v_{1}^{\prime},\ldots,v_{m}^{\prime}$ and $\tau(x)=v_{0}$, DF-QuAD computes $x$’s strength as $\sigma(x)\!=\!\mathcal{C}(v_{0},\mathcal{F}(v_{1},\ldots,v_{n})$, $\mathcal{F}(v_{1}^{\prime},\ldots,v_{m}^{\prime}))$, where, for any $w_{1},\ldots,w_{k}$, $\mathcal{F}(w_{1},\ldots,w_{k})$ is 0 if $k\!=\!0$ and $1-\prod_{i=1}^{k}(|1-w_{i}|)$ otherwise and $\mathcal{C}$ is defined as follows: for $v_{a}=\mathcal{F}(v_{1},\ldots,v_{n})$ and $v_{s}=\mathcal{F}(v_{1}^{\prime},\ldots,v_{m}^{\prime})$, if $v_{a}=v_{s}$ then $\mathcal{C}(v_{0},v_{a},v_{s})=v_{0}$; else if $v_{a}>v_{s}$ then $\mathcal{C}(v_{0},v_{a},v_{s})=v_{0}-(v_{0}\cdot|v_{s}-v_{a}|)$; otherwise $\mathcal{C}(v_{0},v_{a},v_{s})=v_{0}+((1-v_{0})\cdot|v_{s}-v_{a}|)$.
@@ 4 Argumentative Coherence of Forecasters
In this section, we define a novel way of identifying incoherent forecasters, based on their argumentative opinions and predictions. First, we define a novel notion of forecasters as QBAFs drawn from ACFs, as follows.
@@
The base score for the forecasting argument represents the confidence level of the forecaster on this argument prior to debating/voting. Higher assignments of this score mean there is stronger credibility for the forecasting argument, while lower assignments mean the forecasting argument is treated with more skepticism, and intermediate values provide a gradual scale between the two. Varying this base score allows us to control how much weight is given to the forecasting argument independently of the influence of supports and attacks. Also, in a forecaster QBAF, an argument’s base score is 0.5 when the forecaster agrees or disagrees with the argument (we use 0.5 as it is the middle point between 0 and 1). Finally, in a forecaster QBAF, we set the base score to 0 for an argument that the user has not voted on, as we do not want the argument with no opinion to have an effect on the forecasting argument’s strength. The following example illustrates considerations for how the base score for forecasting arguments can be assigned.
@@
Consider a forecasting argument $f$="X will win their next professional tennis match". We may assume a neutral assignment to the forecasting argument (e.g. $\tau(f)=0.5$) as we cannot make any assumptions on the initial confidence. If we know X is an amateur tennis player, then we would expect a lower confidence level, hence a lower assignment to the forecasting argument (e.g. $\tau(f)=0.03$). If instead we know X is the top-ranked player this season, then we would expect a higher confidence level, hence a higher assignment to the forecasting argument (e.g. $\tau(f)=0.7$).
Our forecaster QBAFs differ from those underpinning the original ArguCast Gorur et al. (2023), specifically by satisfying properties of neutral forecasting (ensuring unbiased defaults when no opinions are expressed) and bias-based forecasting (ensuring opinions properly influence prediction strength), which Gorur et al. (2023) fails to satisfy (see Appendix A for details).
@@
We introduce a notion to determine whether a forecaster is argumentatively coherent, which evaluates whether a forecasters’ prediction aligns with their internal reasoning. In particular, we assess the consistency between the forecasters’ opinion (represented by the final strength of the forecasting argument within their QBAF), and their prediction, given two threshold values. This is with the intention of filtering out users who do not satisfy this condition, to potentially gain more accurate forecasts.
@@
In the experiments, we use $\xi_{2}(f)=0.5$ or the initial probabilities from other judgmental forecasting platforms.
@@ 5 In-Person Experiments
In this section we detail the findings from our first user study, which we undertook with members of the public attending [DETAILS OF EVENT OMITTED FOR ANONYMITY]. The participants used the ArguCast system with pre-prepared questions on sports, politics and finance. To assess the usefulness of coherence, we then compared the outcomes of our variant of ArguCast (from now on referred to as ArguCast for simplicity) with those of baseline systems (seen as the ground truth here, considering the much larger basis of forecasters therein, compared to ArguCast).
@@ 5.1 Research Question
We expected that the null hypothesis for RQ1 (i.e., a negative answer to the question) would be rejected.
@@ 5.2 Experimental Design
To explore RQ1, we asked participants to generate debates, express their opinions on the debates, and make predictions (in ArguCast) about three future events, along the following lines: ‘Will England win Euro 2024?’ (Euro), ‘Will the Labour party get a 50% majority on the 4th of July 2024 UK general election?’ (Labour), and ‘Will Nvidia become the most profitable company in the world before the end of September 2024?’ (Nvidia). Participants were given electronic devices such as tablets, laptops, and displays to use ArguCast. The generated debates are illustrated in Appendix B, and the forecaster QBAFs and predictions are available in the supplementary material. Then, we applied our notion of coherence with $\xi_{1}(f)=0.5$ and two options for $\xi_{2}$: $\xi_{2}(f)=0.5$, which is a neutral assignment, and $\xi_{2}(f)=\gamma$, where $\gamma$ is the initial prediction, which we obtained from prediction markets and Metaculus [^5] Specifically, the initial predictions for the Euro and Labour questions were taken from implied probabilities given by www.betfair.com, and the baseline for the Nvidia question was taken from the prediction for this event on www.metaculus.com on the 19th of June 2024.. We then computed the simple average of both the raw predictions (all predictions made by users) and coherent predictions (the predictions not filtered-out by coherence), and compared these predictions with the initial predictions by the baselines to assess the effect of the different values of $\xi_{2}$.
@@ 5.3 Participant Recruitment and Ethics Statement
We recruited 169 anonymous English-speaking participants at [DETAILS OF EVENT OMITTED FOR ANONYMITY]. [^6] Some participants focused only on some of the questions, see Table 1. The data collection was anonymous as we created anonymous accounts for the participants on ArguCast. After reading the participant information sheet, participants who used the ArguCast system gave implied consent. Ethics approval was obtained from our institution prior to conducting the experiments.
@@ 5.4 Results
The results are shown in Table 1. For the Euro question, the raw average prediction (38.06) was far from the initial prediction (20). Applying coherence with $\xi_{2}(f)=0.5$ slightly moved the prediction further from the initial prediction, increasing it to (39.03). In contrast, coherence with $\xi_{2}(f)=\gamma$ brought the prediction down to 34.89, closer to the initial prediction. These results were achieved by filtering out 52 and 56 participants for $\xi_{2}(f)=0.5$ and $\xi_{2}(f)=\gamma$, respectively. For the Labour question, we see a similar trend where the raw average prediction (60.69) was far from the initial prediction (95). Applying coherence improved the results, giving 61.40 with $\xi_{2}(f)=0.5$ and 63.33 with $\xi_{2}(f)=\gamma$. It should be noted that achieving these corrections required filtering out more than half of the participants: out of 54, 29 for $\xi_{2}(f)=0.5$ and 48 for $\xi_{2}(f)=\gamma$, indicating that a substantial proportion of incoherent responses diverged from the initial prediction. For the Nvidia question, the raw average prediction (59.84) was lower than the initial prediction (80). Applying coherence improved on the raw average predictions, 65.10 with $\xi_{2}(f)=0.5$ and 77.00 with $\xi_{2}(f)=\gamma$. Again we filtered out more than half of the participants, 18 for $\xi_{2}(f)=0.5$ and 6 for $\xi_{2}(f)=\gamma$.
@@
Table 1: Predictions for the three forecasting questions comparing the initial prediction (by prediction markets or Metaculus), with the raw average of participant predictions on ArguCast, and the coherent predictions with $\xi_{2}(f)=0.5$ or $\xi_{2}(f)=\gamma$. $N$ is the number of users that participated in the questions.
These results indicate that filtering out incoherent predictions can improve aggregate forecasts. Overall, across all three questions, applying coherence with $\xi_{2}(f)=\gamma$ to ArguCast consistently moved predictions closer to the initial predictions. However, when coherence with $\xi_{2}(f)=0.5$ was applied the prediction moved further from the initial prediction for the Euro question. This suggests that using a neutral coherence threshold can still allow some of the incorrect predictions to be seen as coherent, thus pointing to the importance of a customisable notion of coherence. The results for the Labour and Nvidia questions further support this interpretation, as applying coherence with $\xi_{2}(f)=\gamma$ led to greater improvements over compared to applying rationality with $\xi_{2}(f)=0.5$. Overall, we can reject the null hypothesis for RQ1, since applying coherence (for approrpiate choices of the parameters improves forecasting accuracy.
@@ 6 LLM Experiments
In this section, we present the findings of an empirical experiment conducted using LLMs on forecasting questions from GJOpen (a public judgmental forecasting platform 777www.gjopen.com).
@@ 6.1 Research Question
We expected that the null hypothesis for RQ2 (i.e., a negative answer to the question) would be rejected.

### v2

@@ 1 Introduction
This pre-print is an extended version of Gorur et al. (2025), including supplementary material.
@@ 2 Related Work
In Judgmental forecasting, there is some work that tries to incorporate coherence, for example by investigating the best ways of eliciting probabilities from humans Wallsten and Budescu (1995), how incentives and training change forecasters’ abilities Chang et al. (2016), the effect of scoring rules on forecasters Zellner et al. (2021), and how principles help overcome overconfidence Arkes (2001). Some existing solutions Karvetski et al. (2013); Fan et al. (2019); Gibson et al. (2021) show that forecasting accuracy improves by weighting judges’ assessments based on their degree of coherence. Further, Thomson et al. (2019) develop a measure of coherence to identify the degree of independence between sets of forecasts, thus improving the aggregated forecasts. Additionally, Ho et al. (2024) introduce a notion of probabilistic coherence that identifies superior forecasters. Finally, there has been some work on coherence constraints in hierarchical forecasting Rangapuram et al. (2021); Panagiotelis et al. (2023); Rangapuram et al. (2023); Olivares et al. (2023). In this paper we do not compare our notion against existing approaches, since they do not operate in our setting.
@@
Existing research Irwin et al. (2022); Gorur et al. (2023); Toni et al. (2023) explores the integration of computational argumentation for eliciting, representing and evaluating the argumentative reasoning of users. Specifically, Irwin et al. (2022) were the first to apply computational argumentation to support judgmental forecasting. They introduced their notion of coherence based on how forecasters update their predictions and their reasoning behind those updates. However, as mentioned in the introduction, their underlying framework exhibits some limitations in that it can only handle binary questions and does not allow the same arguments to be used for different questions. It is also important to note that our notion of coherence is based solely on forecasters’ argumentative reasoning and their predictions rather than prediction updates. Further demonstrating the benefits of argumentation in this context is Toni et al. (2023), which incorporates users’ opinions and probabilistic estimates to aid forecasting, but without considering coherence constraints. In addition, the role of argumentation for forecasting is indirectly advocated by resources such as the FORECAST2023 dataset Gorska et al. (2024), where the rationales from the Hybrid-Forecasting Competition are annotated with support/attack relations.
@@
Recent research demonstrates the potential of LLMs in forecasting, e.g. in time-series forecasting, LLMs have been shown to improve accuracy Jin et al. (2024); Tang et al. (2024); Liu et al. (2024). However, direct use of LLMs in judgmental forecasting has yielded mixed results: while some studies Schoenegger and Park (2023); Abolghasemi et al. (2023) show that using these models directly does not improve forecasting accuracy, others Halawi et al. (2024) have concluded that the current LLMs are nearly as good as human forecasters if they are supported with Retrieval Augmented Generation (RAG) Zhao et al. (2024). They further point that the aggregation of forecasts from humans and LLMs with RAG achieve the best forecasting accuracy. Combining human insights with LLM-supported advice has also proven effective in accuracy Schoenegger et al. (2024). Our approach in Section 6 incorporates LLMs on forecaster data to improve forecasting accuracy. It does so by adapting, towards forecasting, ArgLLMs Freedman et al. (2025), an existing system for mining argumentation frameworks from LLMs to verify claims.
@@ 3 Background
QBAFs are equipped with so-called gradual semantics, assessing the (dialectical) strength of arguments. We use the Discontinuity-Free Quantitative Argumentation Debate (DF-QuAD) gradual semantics Rago et al. (2016) to obtain the strength of the forecasters’ opinions. For a given QBAF $\langle\mathcal{X},\mathcal{A},\mathcal{S},\tau\rangle$, for any $x\in\mathcal{X}$ with $n\geq 0$ attackers with strengths $v_{1},\ldots,v_{n}$, $m\geq 0$ supporters with strengths $v_{1}^{\prime},\ldots,v_{m}^{\prime}$ and $\tau(x)=v_{0}$, DF-QuAD computes $x$’s strength as $\sigma(x)\!=\!\mathcal{C}(v_{0},\mathcal{F}(v_{1},\ldots,v_{n})$, $\mathcal{F}(v_{1}^{\prime},\ldots,v_{m}^{\prime}))$, where, for any $w_{1},\ldots,w_{k}$, $\mathcal{F}(w_{1},\ldots,w_{k})$ is 0 if k=0k\!=\!0 and $1-\prod_{i=1}^{k}(|1-w_{i}|)$ otherwise and $\mathcal{C}$ is defined as follows: for $v_{a}=\mathcal{F}(v_{1},\ldots,v_{n})$ and $v_{s}=\mathcal{F}(v_{1}^{\prime},\ldots,v_{m}^{\prime})$, if $v_{a}=v_{s}$ then $\mathcal{C}(v_{0},v_{a},v_{s})=v_{0}$; else if $v_{a}>v_{s}$ then $\mathcal{C}(v_{0},v_{a},v_{s})=v_{0}-(v_{0}\cdot|v_{s}-v_{a}|)$; otherwise $\mathcal{C}(v_{0},v_{a},v_{s})=v_{0}+((1-v_{0})\cdot|v_{s}-v_{a}|)$.
@@ 4 Argumentative Coherence of Forecasters
In this section, we define a novel way of identifying incoherent forecasters, based on their argumentative opinions and predictions. First, we define a novel notion of forecasters as QBAFs drawn from ACFs:
@@
The base score for the forecasting argument represents the confidence level of the forecaster on this argument prior to debating/voting. We choose the [0,1] scale for base scores in-line with existing work in gradual semantics Baroni et al. (2019). Higher assignments of this score mean there is stronger credibility for the forecasting argument, while lower assignments mean the forecasting argument is treated with more skepticism, and intermediate values provide a gradual scale between the two. Varying this base score allows us to control how much weight is given to the forecasting argument independently of the influence of supports and attacks. Also, in a forecaster QBAF, an argument’s base score is 0.5 when the forecaster agrees or disagrees with the argument (we use 0.5 as it is the middle point between 0 and 1). Finally, in a forecaster QBAF, we set the base score to 0 for an argument that the user has not voted on, as we do not want the argument with no opinion to have an effect on the forecasting argument’s strength. The following example illustrates considerations for how the base score for forecasting arguments can be assigned.
@@
Consider a forecasting argument ff="X will win their next professional tennis match". We may assume a neutral assignment to the forecasting argument (e.g. $\tau(f)=0.5$) as we cannot make any assumptions on the initial confidence. If we know X is an amateur tennis player, then we would expect a lower confidence level, hence a lower assignment to the forecasting argument (e.g. $\tau(f)=0.03$). If instead we know X is the top-ranked player this season, then we would expect a higher confidence level, hence a higher assignment to the forecasting argument (e.g. $\tau(f)=0.7$). Thus, the base score reflects facts known to the forecaster prior to debate.
Our forecaster QBAFs differ from those underpinning the original ArguCast Gorur et al. (2023), specifically by satisfying properties of neutral forecasting (ensuring unbiased defaults when no opinions are expressed) and bias-based forecasting (ensuring opinions properly influence prediction strength), which Gorur et al. (2023) fails to satisfy (see Appendix A).
@@
We introduce a notion to determine whether a forecaster is argumentatively coherent. This notion constrains a forecasters’ prediction to align with their internal reasoning. In particular, we assess the consistency between the forecasters’ opinion (represented by the final strength of the forecasting argument within their QBAF), and their prediction, given two threshold values. This is with the intention of filtering out users who do not satisfy this condition, to potentially gain more accurate forecasts.
@@
Both threshold values are task-specific and thus need to be optimised for each task. In the experiments, we use $\xi_{2}(f)=0.5$ or the initial probabilities from other judgmental forecasting platforms.
@@ 5 In-Person Experiments
In this section we detail the findings from our first user study, which we undertook with members of the public attending the Great Exhibition Road Festival (GERF) in London in June 2024555www.greatexhibitionroadfestival.co.uk. The participants used the ArguCast system with pre-prepared questions on sports, politics and finance. To assess the usefulness of coherence, we then compared the outcomes of our variant of ArguCast (from now on referred to as ArguCast for simplicity) with those of baseline systems (seen as the ground truth here, considering the much larger basis of forecasters therein, compared to ArguCast).
@@ 5.1 Research Question
We expected the null hypothesis (i.e., a negative answer to the question) to be rejected.
@@ 5.2 Experimental Design
To explore RQ1, we asked participants to generate debates, express their opinions on the debates, and make predictions (in ArguCast) about three future events, along the following lines: ‘Will England win Euro 2024?’ (Euro), ‘Will the Labour party get a 50% majority on the 4th of July 2024 UK general election?’ (Labour), and ‘Will Nvidia become the most profitable company in the world before the end of September 2024?’ (Nvidia). Participants were given electronic devices such as tablets, laptops, and displays to use ArguCast. The generated debates are illustrated in Appendix B, and the forecaster QBAFs and predictions are available in the supplementary material. Then, we applied our notion of coherence with $\xi_{1}(f)=0.5$ and two options for $\xi_{2}$: $\xi_{2}(f)=0.5$, which is a neutral assignment, and $\xi_{2}(f)=\gamma$, where $\gamma$ is the initial prediction, which we obtained from prediction markets and Metaculus [^6] Specifically, the initial predictions for the Euro and Labour questions were taken from implied probabilities given by www.betfair.com, and the baseline for the Nvidia question was taken from the prediction for this event on www.metaculus.com on the 19th of June 2024.. We then computed the simple average of both the raw predictions (made by users) and coherent predictions (the predictions not filtered-out by coherence), and compared these predictions with the initial predictions by the baselines to assess the effect of the different values of $\xi_{2}$.
@@ 5.3 Participant Recruitment and Ethics Statement
We recruited 169 anonymous English-speaking participants at the GERF. [^7] Some participants focused only on some of the questions, see Table 1. The data collection was anonymous as we created anonymous accounts for the participants on ArguCast. After reading the participant information sheet, participants who used the ArguCast system gave implied consent. Ethics approval was obtained from our institution prior to conducting the experiments.
@@ 5.4 Results
The results are shown in Table 1.888Due to the small sample size and exploratory design, we did not perform statistical testing. For the Euro question, the raw average prediction (38.06) was far from the initial prediction (20). Applying coherence with $\xi_{2}(f)\!=\!0.5$ slightly moved the prediction further from the initial prediction, increasing it to (39.03). In contrast, coherence with $\xi_{2}(f)\!=\!\gamma$ brought the prediction down to 34.89, closer to the initial prediction. These results were achieved by filtering out 52 and 56 participants for $\xi_{2}(f)\!=\!0.5$ and $\xi_{2}(f)\!=\!\gamma$, respectively. For the Labour question, we see a similar trend where the raw average prediction (60.69) was far from the initial prediction (95). Applying coherence improved the results, giving 61.40 with $\xi_{2}(f)\!=\!0.5$ and 63.33 with $\xi_{2}(f)\!=\!\gamma$. It should be noted that achieving these corrections required filtering out more than half of the participants: out of 54, 29 for $\xi_{2}(f)\!=\!0.5$ and 48 for $\xi_{2}(f)\!=\!\gamma$, indicating that a substantial proportion of incoherent responses diverged from the initial prediction. For the Nvidia question, the raw average prediction (59.84) was lower than the initial prediction (80). Applying coherence improved on the raw average predictions, 65.10 with $\xi_{2}(f)\!=\!0.5$ and 77.00 with $\xi_{2}(f)\!=\!\gamma$. Again we filtered out more than half of the participants, 18 for $\xi_{2}(f)\!=\!0.5$ and 6 for $\xi_{2}(f)\!=\!\gamma$.
@@
Table 1: Predictions for the three forecasting questions comparing the initial prediction (by prediction markets or Metaculus), with the raw average of participant predictions on ArguCast, and the coherent predictions with $\xi_{2}(f)\!=\!0.5$ or $\xi_{2}(f)\!=\!\gamma$. NN is the number of participating users.
These results indicate that filtering out incoherent predictions can improve aggregate forecasts. Overall, across all three questions, applying coherence with $\xi_{2}(f)=\gamma$ to ArguCast consistently moved predictions closer to the initial predictions. However, when coherence with $\xi_{2}(f)=0.5$ was applied the prediction moved further from the initial prediction for the Euro question. This suggests that using a neutral coherence threshold can still allow some of the incorrect predictions to be seen as coherent, thus pointing to the importance of a customisable notion of coherence. The results for the Labour and Nvidia questions further support this interpretation, as applying coherence with $\xi_{2}(f)=\gamma$ led to greater improvements over compared to applying rationality with $\xi_{2}(f)=0.5$. The consistency in these improvements is valuable, particularly given the complexity and noise inherent in human judgment data Kahneman (2011). Overall, we can reject the null hypothesis for RQ1, since applying coherence (for appropriate choices of the parameters improves forecasting accuracy.
@@ 6 LLM Experiments
In this section, we present the findings of an empirical experiment conducted using LLMs on forecasting questions from GJOpen (a public judgmental forecasting platform 999www.gjopen.com).
@@ 6.1 Research Question
We expected the null hypothesis (i.e., a negative answer to the question) to be rejected.